A hardworking ant must supply 0.00655 N to pull a small piece of fruit at constant velocity over a distance of 9.40 cm up a small hill to the entrance of the colony. lIf the coefficient of kinetic friction between the piece of fruit and the 20.8 degree sloped hill is 0.540, calculate the work done by the ant by pulling the piece of fruit up the hill. Use the work-energy theorem to calculate the mass of the piece of fruit in grams. (The acceleration due to gravity is 9.81 m/s2).
Answer #1

frictional force = Uk*mg*cos(20.8 degrees) = 0.540*m*9.8*0.934 N

gravitational force along the incline = mgsin(20.8 degrees) = m*9.8*0.355 N

moving at constant velcity => net acceleration = 0

=> 0.00655 = 0.540*m*9.8*0.934 + m*9.8*0.355

=> mass of fruit m = 0.777 gram

and wordone = F*S = 0.00655*9.40 = 0.0615 J